find the quadratic function whose vertex is (-1,5) and passing through the point (2,7) and graph the function.
need answer before 12/1/08

Answer by scott8148(6628)

using the vertex form of the equation __ y=a(x+1)^2+5

substituting the point to find a __ 7=a(2+1)^2+5

subtracting 5 __ 2=9a

dividing by 9 __ 2/9=a

y=(2/9)(x+1)^2+5 __ y=(2/9)x^2+(4/9)x+(47/9)